### v1

# Dual Target-Mounted RISs-Assisted ISAC Against Eavesdropping and Malicious Interference

Abstract
The synergy between integrated sensing and communication (ISAC) and reconfigurable intelligent surfaces (RISs) unlocks novel applications and advanced services for next-generation wireless networks, yet also introduces new security challenges. In this study, a novel dual target-mounted RISs-assisted ISAC scheme is proposed, where a base station with ISAC capability performs sensing of two unmanned aerial vehicle (UAV) targets, one of which is legitimate and the other is eavesdropper, while communicating with the users through an RIS mounted on the legitimate UAV target. The proposed scheme addresses dual security threats posed by a hostile UAV target: eavesdropping on legitimate user communications and random interference attacks launched by a malicious RIS mounted on this eavesdropper UAV target, aiming to disrupt secure transmissions. A non-convex optimization problem maximizing the secrecy rate of the users is formulated, and a semi-definite relaxation (SDR)-based two-stage solution is developed to optimize the transmit beamforming matrix of the base station and the phase shift coefficients of the legitimate RIS. Extensive computer simulations are conducted to evaluate the robustness of the proposed solution under various system configurations. The proposed system’s communication performance is assessed using the secrecy rate metric, while the sensing performance is evaluated through the signal-to-interference-plus-noise ratio and the Cramer–Rao bound (CRB) for angle-of-departure (AoD) estimation of the eavesdropper UAV target.
Index Terms:
Integrated sensing and communication (ISAC), reconfigurable intelligent surface (RIS), secure communication, semi-definite relaxation (SDR).
This work was supported by The Scientific and Technological Research Council of Turkiye (TUBITAK) through the $15$ Frontier Research and Development Laboratories Support Program under Project 5229901 - 6GEN. Lab: 6G and Artificial Intelligence Laboratory, and also Grant 120E401.
I Introduction
The paradigm shift from supporting communication solely between people to seamless interactions between people and connected devices has laid a foundation for a new generation demanding enhanced data rates, ultra-low latency, and massive device connectivity [1]. As the sixth-generation (6G) era approaches, the foundational framework of its evolution has begun to shape through International Telecommunication Union (ITU) recommendations [2] and ongoing 3rd Generation Partnership Project (3GPP) releases [3]. Accordingly, while building upon enhanced fifth-generation capabilities, 6G introduces three novel transformative usage scenarios: artificial intelligence-driven communication, ubiquitous connectivity, and integrated sensing and communication (ISAC), supporting diverse emerging applications from autonomous systems, satellite-connected networks, and advanced environmental perception [2].
In 6G networks, ISAC technology offers enabling communication and environmental sensing in a unified hardware platform, exploiting shared frequency spectrum, resources, and signal processing capabilities [4]. This opens up a wide range of applications across various environment-aware scenarios, including environmental monitoring, smart cities, autonomous transportation, and internet of things (IoT) [4]. These applications leverage newly introduced advanced sensing-related capabilities such as high-precision mapping, sub-centimeter resolution imaging, and centimeter-level localization enabled by ISAC technologies in 6G networks [2].
Although ISAC is now widely recognized as a key 6G technology, its earliest studies trace back to several decades, exploring the dual use of radar and communication systems [4]. Since then, this research has appeared under various terminologies, such as joint radar and communication (JRC) [5], joint communication and sensing (JCAS) [6], and dual-function radar-communication (DFRC) [7], ultimately converging towards the concept of ISAC [4].
Another promising candidate technology in 6G networks is reconfigurable intelligent surfaces (RISs)-empowered communication, offering a controllable wireless environment in a cost-effective manner [8]. RISs are fundamentally based on meta-surface technology employing tunable electronic components such as PIN diodes or varactor diodes to dynamically control electromagnetic (EM) properties of incident signals [9, 8]. By switching between different states of these elements, RISs can manipulate EM characteristics of incident waves such as phase, amplitude, and polarization [8]. This unprecedented capability of RISs positions them as a strong candidate technology for 6G wireless networks [9]. Over the past few years, research on RIS has primarily concentrated on improved signal quality [9], physical layer security [10], enhanced capacity achievement [11], and passive beamforming designs [12], which leverages RIS to shape the propagation environment in a constructive manner. However, a few studies have also examined RIS as a potential security threat, demonstrating how maliciously configured surfaces can degrade a legitimate wireless system [13, 14].
As 6G studies advance, a growing number of studies for RIS-assisted ISAC schemes have emerged in the literature [15]. In [16, 17], the sensing-capable RIS concepts have been introduced to tackle the inherent severe path attenuation of RIS-aided system designs. In [18, 19], a joint beamforming and phase optimization algorithm has been developed to satisfy quality of service (QoS) requirements of both communication users and sensing targets. Moreover, target-mounted aerial RIS-aided ISAC schemes have been presented to improve coverage [20] and facilitate accurate localization [16]. Furthermore, several studies have explored secure communication in RIS-assisted ISAC systems, considering scenarios where sensing targets could act as potential eavesdroppers intercepting legitimate users, while RISs are strategically configured to overcome security challenges effectively [21, 22]. Although prior studies have investigated the effects of eavesdropper targets in RIS-aided ISAC schemes [20, 21, 22, 23, 24], the joint impact of eavesdropper targets and malicious RIS attacks within an ISAC system has not yet been investigated.
In this paper, to overcome the above challenges, a novel RIS-assisted ISAC system is proposed, where a base station (BS) equipped with a uniform linear array (ULA) performs sensing of two unmanned aerial vehicle (UAV) targets while simultaneously establishing reliable communication links with the users through an RIS mounted on the legitimate UAV. The system model considers security threats posed by an eavesdropper UAV target that intercepts wireless communication signals while a malicious RIS, mounted on the eavesdropper UAV, attempts to disrupt the users’ communication by launching random interference attacks. To investigate this dual-threat scenario, including both an eavesdropper UAV and a malicious RIS, a semi-definite relaxation (SDR)-based two-stage optimization algorithm is developed to determine the transmit beamforming matrix of the BS and the reflection coefficients of the legitimate RIS such that the system secrecy rate is maximized. The system performance is evaluated through computer simulations by analyzing the communication subsystem with the users’ secrecy rate and the sensing subsystem with the sensing signal-to-interference ratio (SINR) of UAV targets and their Cramer-Rao bounds (CRBs) for the two-dimensional (2D) angle of departures (AoDs) estimations. The major contributions of this study can be summarized as follows:
  • A dual target-mounted RISs-assisted secure ISAC system is considered, where one UAV target both acts as an eavesdropper to intercept the communication between the BS and the users and launches random interference attacks using the malicious RIS to disrupt the user communication links. To remedy the eavesdropper attacks, an RIS mounted on the legitimate UAV is utilized to enhance the system secrecy rate while satisfying the sensing performance requirements.
  • The optimization problem for the proposed system model is formulated to maximize the system secrecy rate while guaranteeing achievable sensing performance under the attacks of both eavesdropping and malicious RIS interference. To address this challenge, an SDR-based two-stage solution is developed to optimize both the transmit beamforming matrix of the BS and the phase shift coefficients of the legitimate RIS.
  • As a communication performance metric, the sum secrecy rate of the legitimate users is investigated through extensive computer simulations. For sensing performance, the SINR of UAV targets is analyzed, and the CRB is derived to estimate the 2D AoDs of UAV targets.
  • The simulation results demonstrate the effectiveness of the proposed algorithm, showing a notable performance improvement in the secrecy rate of the communication users while simultaneously enhancing key sensing performance metrics, namely SINR and CRB.
The rest of the paper is organized as follows. Section II introduces the proposed UAV-mounted RIS-assisted secure ISAC framework. In Section III, the problem formulation for beamforming design and legitimate RIS phase optimization is provided. Section IV provides the performance evaluation for both the communication and sensing subsystems, and Section V concludes the paper.
Notation: Unless otherwise specified, scalars are denoted by italic letters (i.e., $x$), while vectors and matrices are represented by boldface lower letters (i.e., $\mathbf{x}$) and boldface upper letters (i.e., $\mathbf{X}$), respectively. $\mathrm{diag}(\mathbf{\mathbf{x}})$ represents a diagonal matrix whose diagonal elements are the elements of the vector $\mathbf{x}$, while $\mathbf{I}$ denotes the identity matrix. $|x|$ and $\left\|\mathbf{X}\right\|_{\mathrm{F}}$ stand for the absolute value of a scalar and the Frobenius norm of a matrix, respectively. $x^{*}$ is the conjugate of $x$, while $\mathbf{X}^{-1}$, $\mathbf{X}^{\mathrm{T}}$, and $\mathbf{X}^{\mathrm{H}}$ represent the inverse, transposition, and Hermitian of a matrix, respectively. $\operatorname{vec}(\cdot)$ stands for vectorization operator, while $\operatorname{Tr}(\cdot)$ represents trace operator. $\otimes$ and $\odot$ stand for Kronecker and Hadamard products, respectively. $\mathfrak{Re}(\cdot)$ and $\mathfrak{Im}(\cdot)$ denote real and imaginary components of a complex number, respectively. $\mathcal{CN}(\varpi,\sigma^{2})$ represents the distribution of a complex Gaussian random variable with mean $\varpi$ and variance $\sigma^{2}$. $\mathbb{C}^{M\times L}$ denotes the space of complex matrices with dimensions of $M\times L$. $\mathcal{O}$ stands for big-O notation and $\mathbf{X}\succeq 0$ denotes positive semi-definiteness of matrix $\mathbf{X}$.
II System Model
In this section, the system model of the proposed UAV-mounted RISs-assisted secure ISAC system is presented.
II-A Dual UAV-mounted RISs-assisted ISAC System
In the proposed scheme, as illustrated in Fig. 1, a BS, equipped with $T_{x}$ transmit antennas in a ULA structure, performs sensing of two UAV targets while simultaneously establishing communication links with $K$ single-antenna legitimate users via RISs mounted on those UAVs. Among two UAV targets, one is legitimate, while the other attempts to eavesdrop on the communication users. The legitimate UAV target carries a legitimate RIS to sustain reliable communication links between the BS and the users, whereas the eavesdropping UAV target is equipped with a malicious RIS, introducing random interference attacks to disrupt the communication links between the BS and the users. Both legitimate and malicious RISs are passive with uniform planar arrays (UPAs), consisting of $N_{L}$ and $N_{M}$ reflecting elements, respectively. The reflection matrices of legitimate and malicious RIS can be respectively given as
Figure 1: Dual Target-mounted RISs-assisted ISAC scheme.
$\mathbf{\Omega}_{L}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathrm{diag}\left(e^{j\psi_{1}},e^{j\psi_{2}},\cdots,e^{j\psi_{N_{L}}}\right)$ (1)
$\mathbf{\Omega}_{M}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathrm{diag}\left(e^{j\zeta_{1}},e^{j\zeta_{2}},\cdots,e^{j\zeta_{N_{M}}}\right)$ (2)
where $\psi_{l}$ is the phase shift of $l$-th reflecting element of the legitimate RIS and $\zeta_{m}$ is the phase shift of the $m$-th reflecting element of the malicious RIS, for $l\in\{1,\cdots,N_{L}\}$ and $m\in\{1,\cdots,N_{M}\}$. Here, since the malicious RIS performs random interference attacks, each $\zeta_{m}$ is assumed to be independently and identically distributed (i.i.d.) complex Gaussian random variable and follows $\sim\mathcal{CN}(0,1)$ distribution, while $\psi_{l}$ is optimized for enhancing secure transmission. In the proposed scheme, for enabling joint sensing and communication simultaneously, the BS transmits a unified signal at the $t$-th time interval as follows
$\mathbf{x}(t)\in\mathbb{C}^{T_{x}\times 1}=\mathbf{W}_{c}{\mathbf{x}_{c}}(t)+\mathbf{W}_{s}{\mathbf{x}_{s}}(t)$ (3)
$\hskip 55.48277pt=\bigl{[}\mathbf{W}_{c},\mathbf{W}_{s}\bigr{]}\bigl{[}{\mathbf{x}_{c}}(t),{\mathbf{x}_{s}}(t)\bigr{]}^{\mathrm{T}}$ (4)
where $\mathbf{W}_{c}\in\mathbb{C}^{T_{x}\times K}$ and $\mathbf{W}_{s}\in\mathbb{C}^{T_{x}\times T_{x}}$ are the transmit beamforming matrix for communication users and sensing UAV targets, respectively. Here, ${\mathbf{x}_{c}}(t)\in\mathbb{C}^{K\times 1}$ is the communication signal for the legitimate users and satisfying $\mathbb{E}\{\mathbf{c}(t)\mathbf{x}_{c}(t)^{\mathrm{H}}\}=\mathbf{I}_{K}$, while and ${\mathbf{x}_{s}}(t)\in\mathbb{C}^{T_{x}\times 1}$ is frequency modulated continuous wave (FMCW) sensing signal [25] satisfying $\mathbb{E}\{\mathbf{x}_{s}(t)\mathbf{x}_{s}(t)^{\mathrm{H}}\}=\mathbf{I}_{T_{x}}$ and also orthogonal to the communication signal to prevent mutual information $\mathbb{E}\{\mathbf{x}_{s}(t)\mathbf{x}_{c}(t)^{\mathrm{H}}\}=\mathbf{0}_{T_{x}\times K}$. Therefore, for $\mathbf{W}\in\mathbb{C}^{T_{x}\times(T_{x}+K)}=\bigl{[}\mathbf{W}_{c},\mathbf{W}_{s}\bigr{]}$ being the overall transmit beamforming matrix, the covariance of the transmit signal becomes
$\mathbf{R_{x}}=\mathbb{E}\{\mathbf{x}(t)\mathbf{x}(t)^{\mathrm{H}}\}$ (5)
$\hskip 14.22636pt=\mathbf{W}\mathbf{W}^{\mathrm{H}}.$ (6)
Therefore, the maximum total transmit power constraint at the BS can be calculated as
$P_{T}\geq\mathrm{Tr}(\mathbf{R}_{x}).$ (7)
II-B Sensing Model
In the proposed scheme, both the legitimate and eavesdropper UAV targets are assumed to have line-of-sight (LoS) links with the BS. The array responses between the BS and the UAV targets are modeled using the 2D AoD, defined by the horizontal and vertical angles $\theta^{\text{BS}}_{i}$ and $\phi^{\text{BS}}_{i}$, respectively, where $i\in\{L,E\}$ indicates legitimate UAV ($L$-UAV) and eavesdropper UAV ($E$-UAV), respectively. Then, the transmit steering vector for $i$-UAV is given by
$\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\in\mathbb{C}^{1\times T_{x}}=[1,\cdots,e^{j\nu_{T}(T_{x}-1)\cos(\theta^{\text{BS}}_{i})\cos(\phi^{\text{BS}}_{i})}]$ (8)
where $\nu_{T}=\frac{2\pi}{\lambda}d_{T}$ for $\lambda$ being the waveform and $d_{T}$ being the spacing between adjacent antenna elements. Therefore, the received echo signal at the BS from $i$-UAV target over $L_{s}$ coherent time block becomes
$\mathbf{Y}_{i}=\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\mathbf{X}+\mathbf{N}_{i}$ (9)
where $\mathbf{X}=[\mathbf{x}(1),\cdots,\mathbf{x}(L_{s})]$ and $\mathbf{N}_{i}\in\mathbb{C}^{T_{x}\times L_{s}}$ is the additive white Gaussian noise (AWGN) whose each entry follows $\mathcal{CN}(0,\sigma^{2}_{n})$ distribution. Here, $\beta_{i}$ is the complex-valued round-trip path attenuation between the BS and $i$-UAV [26], which can be given as
$\beta_{i}=\sqrt{\frac{\lambda^{2}S}{64\pi^{3}d_{i}^{4}}}$ (10)
where $S$ is the radar cross section (RCS). Therefore, the sensing SINR of the $i$-UAV target can be calculated as
$\gamma_{i}=\frac{\left\|\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\mathbf{W}\right\|_{\mathrm{F}}^{2}}{\left\|\beta_{j}\mathbf{a}(\theta^{\text{BS}}_{j},\phi^{\text{BS}}_{j})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{j},\phi^{\text{BS}}_{j})\mathbf{W}\right\|_{\mathrm{F}}^{2}+\sigma_{n}^{2}}$ (11)
where $j\neq i\in\{L,E\}$.
For further sensing performance evaluation, the CRB estimation of AoD pairs of legitimate and eavesdropper UAV targets is derived from the Fisher information matrix (FIM). In order to determine FIM of the $i$-UAV, first, the received echo signal given in (9) is vectorized as
$\mathbf{y}_{i}=\mathbf{p}_{i}+\mathbf{v}_{i}$ (12)
where $\mathbf{p}_{i}=\operatorname{vec}(\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\mathbf{X})$ and $\mathbf{v}_{i}=\operatorname{vec}(\mathbf{N}_{i})$. Then, for $\boldsymbol{\omega}_{i}\in\mathbb{C}^{4\times 1}=[\boldsymbol{\vartheta}_{i},\boldsymbol{\delta}_{i}]$, and $\boldsymbol{\vartheta}_{i}=[\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i}]$ and $\boldsymbol{\delta}_{i}=[\mathfrak{Re}(\beta_{i}),\mathfrak{Im}(\beta_{i})]^{\mathrm{T}}$ the FIM for $i$-UAV $\mathbf{F}_{i}\in\mathbb{C}^{4\times 4}$[27] can be expressed as follows
$\mathbf{F}_{i}=\begin{bmatrix}\mathbf{F}_{\boldsymbol{\vartheta}_{i}\boldsymbol{\vartheta}_{i}}&\mathbf{F}_{\boldsymbol{\vartheta}_{i}\boldsymbol{\delta}_{i}}\\ \mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\vartheta}_{i}}&\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\delta}_{i}}\end{bmatrix}$ (13)
while the CRB estimation of AoD pairs of the $i$-UAV target is calculated as
$\text{CRB}(\boldsymbol{\vartheta}_{i})=[\mathbf{F}_{\boldsymbol{\vartheta}_{i}\boldsymbol{\vartheta}_{i}}-\mathbf{F}_{\boldsymbol{\vartheta}_{i}\boldsymbol{\delta}_{i}}(\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\delta}_{i}})^{-1}\mathbf{F}_{\boldsymbol{\boldsymbol{\delta}_{i}\vartheta}_{i}}]^{-1}.$ (14)
Here, each element of the FIM matrix in (13) can be given as
$\mathbf{F}_{i}(\epsilon,\tau)=\frac{2}{\sigma_{n}^{2}}\mathfrak{Re}\left(\frac{{d}\mathbf{p}_{i}^{\mathrm{H}}}{d\boldsymbol{\omega}_{i,\epsilon}}\frac{{d}\mathbf{p}_{i}}{d\boldsymbol{\omega}_{i,\tau}}\right)$ (15)
where $\boldsymbol{\omega}_{i,\epsilon}$ and $\boldsymbol{\omega}_{i,\tau}$ are the $\epsilon$- and $\tau$-th elements of $\boldsymbol{\omega}_{i}$ for ${\epsilon,\tau}\in\{1,2,3,4\}$. Considering the transmit steering vector of $i$-UAV target $\mathbf{a}(\theta_{i}^{\text{BS}},\phi_{i}^{\text{BS}})$ given in (8), let $\mathbf{A}_{i}=\mathbf{a}^{\mathrm{H}}(\theta_{i}^{\text{BS}}\phi_{i}^{\text{BS}})\mathbf{a}(\theta_{i}^{\text{BS}}\phi_{i}^{\text{BS}})$, while $\bar{\mathbf{A}}_{\theta_{i}^{\text{BS}}}$, $\bar{\mathbf{A}}_{\phi_{i}^{\text{BS}}}$ and $\bar{\mathbf{A}}_{\boldsymbol{\delta}_{i}}$ be the derivatives of $\mathbf{A}_{i}$ with respect to $\theta_{i}^{\text{BS}}$, $\phi_{i}^{\text{BS}}$ and $\boldsymbol{\delta}_{i}$}, respectively. Therefore, the FIM matrix elements become [27, 28]
$\mathbf{F}_{\boldsymbol{\vartheta}_{i}\boldsymbol{\vartheta}_{i}}=$
[…]
$\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\vartheta}_{i}}=\frac{2L_{s}}{\sigma_{n}^{2}}\mathfrak{Re}\left(\begin{bmatrix}{\beta_{i}^{*}\operatorname{Tr}(\mathbf{A}}_{i}\mathbf{R_{x}}(\bar{\mathbf{A}}_{\theta_{i}^{\text{BS}}})^{\mathrm{H}})\\ {\beta_{i}^{*}\operatorname{Tr}(\mathbf{A}}_{i}\mathbf{R_{x}}(\bar{\mathbf{A}}_{\phi_{i}^{\text{BS}}})^{\mathrm{H}})\end{bmatrix}[1,j]\right)$ (17)
$\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\delta}_{i}}=\frac{2L_{s}}{\sigma_{n}^{2}}\mathbf{I}_{2}\mathbf{R_{x}}{\operatorname{Tr}(\mathbf{A}}_{i}\mathbf{R_{x}}(\bar{\mathbf{A}}_{i})^{\mathrm{H}}).$ (18)
II-C Communication Model
In the proposed scheme, the communication channels between the BS and RISs, as well as between the RISs and users, are modeled using Rician fading. Let $\mathbf{H}_{n}\in\mathbb{C}^{N_{n}\times T_{x}}$ denote the channel matrix between the BS and the $n$-th RIS, where $n\in\{L,M\}$ represents the legitimate ($L$-RIS) or malicious ($M$-RIS) surface. Similarly, let $\mathbf{g}_{n,k}\in\mathbb{C}^{1\times N_{n}}$ denote the channel vector between the $n$-th RIS and single-antenna user $k$ (Uk) , for $k\in\{1,\cdots,K\}$. These channels can be expressed as
$\mathbf{H}_{n}=\sqrt{\frac{{L_{0}}}{d_{n}^{\alpha_{n}}}}\left({\sqrt{\frac{\kappa}{1+\kappa}}\mathbf{H}^{\text{LOS}}_{n}+\sqrt{\frac{1}{1+\kappa}}\mathbf{H}^{\text{NLOS}}_{n}}\right)$ (19)
$\mathbf{g}_{n,k}=\sqrt{\frac{{L_{0}}}{d_{n,k}^{\alpha_{k}}}}\left(\sqrt{\frac{\kappa}{1+\kappa}}\mathbf{g}_{n,k}^{\text{LOS}}+\sqrt{\frac{1}{1+\kappa}}\mathbf{g}_{n,k}^{\text{NLOS}}\right)$ (20)
where $\kappa$ denotes the Rician factor, $L_{0}$ is the reference path loss at a distance of $1$ meter (m), and $d_{n}$ and $d_{n,k}$ represent the distances between the BS and the $n$-RIS, and between the $n$-RIS and Uk, respectively. The corresponding path loss exponents are denoted by $\alpha_{n}$ and $\alpha_{n,k}$ for the BS–$n$-RIS and $n$-RIS–Uk links, respectively. In the proposed scheme, each element of the NLOS components of the communication channels $\mathbf{H}^{\text{NLOS}}_{n}$ and $\mathbf{g}^{\text{NLOS}}_{n,k}$ is assumed to be i.i.d. and following $\mathcal{CN}(0,1)$ distribution. On the other hand, $\mathbf{H}^{\text{LOS}}_{n}$ and $\mathbf{g}^{\text{LOS}}_{n,k}$ LOS components are deterministic and generated by steering vectors as follows:
$\mathbf{H}^{\text{LOS}}_{n}=\mathbf{b}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})^{\mathrm{H}}\mathbf{b}(\theta_{n}^{\text{BS}},\phi_{n}^{\text{BS}})$ (21)
$\mathbf{g}^{\text{LOS}}_{n,k}=\mathbf{b}(\theta_{n,k}^{\text{RIS}},\phi_{n,k}^{\text{RIS}})$ (22)
where $\mathbf{b}(\theta_{n}^{\text{BS}},\phi_{n}^{\text{BS}})\in\mathbb{C}^{1\times T_{x}}$ is the ULA steering vector of the BS for the corresponding $\{\theta_{n}^{\text{BS}},\phi_{n}^{\text{BS}}\}$ horizontal and vertical parts of the 2D AoD, while the UPA steering vectors of $n$-RIS towards BS and Uk are represented by $\mathbf{b}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\in\mathbb{C}^{1\times N_{n}}$ and $\mathbf{b}(\theta_{n,k}^{\text{RIS}},\phi_{n,k}^{\text{RIS}})\in\mathbb{C}^{1\times N_{n}}$, for the AoD of $\{\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}}\}$ and $\{\theta_{n,k}^{\text{RIS}},\phi_{n,k}^{\text{RIS}}\}$, respectively. Therefore, they can be given as
$\mathbf{b}(\theta_{n}^{\text{BS}},\phi_{n}^{\text{BS}})=[1,\cdots,e^{j{\nu_{T}}(T_{x}-1)\cos(\theta^{\text{BS}}_{n})\cos(\phi^{\text{BS}}_{n})}]$ (23)
$\mathbf{b}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})=\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\otimes\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})$ (24)
where $\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\in\mathbb{C}^{1\times N_{n}^{x}}$ and $\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\in\mathbb{C}^{1\times N_{n}^{z}}$ represent the steering vectors of the $n$-RIS along the $x$- and $z$-axes, respectively. Here, $N_{x}^{n}$ and $N_{z}^{n}$ denote the number of reflecting elements along the $x$- and $z$-directions, and the total number of elements is given by $N_{n}=N_{x}^{n}\times N_{z}^{n}$ for $n\in\{L,M\}$. Therefore, $\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})$ and $\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})$ can be given as
$\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})=[1,\cdots,e^{j{\nu_{R}}(N_{n}^{x}-1)\cos(\theta^{\text{RIS}}_{n})\cos(\phi^{\text{RIS}}_{n})}]$ (25a)
$\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})=[1,\cdots,e^{j{\nu_{R}}(N_{n}^{z}-1)\sin(\phi^{\text{RIS}}_{n})}]$ (25b)
where $\{\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}}\}$ is the AoD of the $n$-RIS towards BS and $\nu_{R}=\frac{2\pi}{\lambda}d_{R}$ for $d_{R}$ being the distance between two horizontally or vertically adjacent RIS elements. Similarly, the LOS component $\mathbf{g}_{n,k}^{\text{LOS}}(\theta_{n,k}^{\text{RIS}},\phi_{n,k}^{\text{RIS}})$ in (22) can be formulated as in (25), based on the AoD pair $\{\theta_{n,k}^{\text{RIS}},\phi_{n,k}^{\text{RIS}}\}$.
The received signal at Uk, incorporating the reflections from both legitimate and malicious RISs, can be given as
$y_{k}(t)=\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{x}(t)+\mathbf{g}_{M,k}\mathbf{\Omega}_{M}\mathbf{H}_{M}\mathbf{x}(t)+n_{k}(t)$ (26)
where $n_{k}$ is the AWGN figure with $\mathcal{CN}(0,\sigma_{n}^{2})$ distribution. Therefore, the SINR of the Uk, given the overall beamforming matrix as $\mathbf{W}\in\mathbb{C}^{T_{x}\times(K+T_{x})}=[\mathbf{w}_{1},\cdots,\mathbf{w}_{K+T_{x}}]$, is expressed as
$\eta_{k}=\frac{\left\|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{k}\right\|_{\mathrm{F}}^{2}}{\sum_{\hat{k}}\left\|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{\hat{k}}\right\|_{\mathrm{F}}^{2}+\left\|\mathbf{g}_{M,k}\mathbf{\Omega}_{M}\mathbf{H}_{M}\mathbf{W}\right\|_{\mathrm{F}}^{2}+\sigma_{n}^{2}}$ (27)
[…]
II-D Security Model
In the proposed scheme, since communication signals through users are assumed to be intercepted by an eavesdropper UAV, using (8), the received signal at the eavesdropper UAV target at the $t$-th time index can be expressed as
$y_{E}(t)=\sqrt{\beta_{E}}\mathbf{a}(\theta_{E}^{\text{BS}},\phi_{E}^{\text{BS}})^{\mathrm{H}}\mathbf{x}(t)+n_{E}(t)$ (28)
where $n_{E}(t)$ is AWGN following $\mathcal{CN}(0,\sigma_{n}^{2})$. Therefore, the eavesdropper SINR on the $k$-th legitimate user can be calculated as
$\eta_{E,k}=\frac{\left\|\mathbf{a}(\theta_{E}^{\text{BS}},\phi_{E}^{\text{BS}})^{\mathrm{H}}\mathbf{w}_{k}\right\|_{\mathrm{F}}^{2}}{\sum_{\hat{k}}\left\|\mathbf{a}(\theta_{E}^{\text{BS}},\phi_{E}^{\text{BS}})^{\mathrm{H}}\mathbf{w}_{\hat{k}}\right\|_{\mathrm{F}}^{2}+\sigma_{n}^{2}}$ (29)
[…]
$S_{R,k}=[R_{k}-R_{E,k}]^{+}$ (30)
where $x]^{+}=\max(x,0)$. Here, $R_{k}=\log_{2}(1+\eta_{k})$ is the achievable rate of user Uk and $R_{E,k}=\log_{2}(1+\eta_{E,k})$ represents the eavesdropping rate at for Uk.
III Problem Formulation and Proposed Solution
In the proposed scheme, the BS operates in dual mode for communication and sensing purposes, yet it is unaware of the malicious RIS mounted on the $E$-UAV. This malicious RIS forms its reflection matrix $\mathbf{\Omega}_{M}$ in (2) using randomly generated phase shifts. To counteract security threats of the $M$-RIS and its mounted $E$-UAV interception threats, the reflection coefficients of the legitimate RIS and the BS transmit beamforming matrix are optimized to maximize the system secrecy rate in (30). To address this, the following optimization problem is formulated:
$\text{(P1)}\hskip 11.38092pt\max_{\mathbf{W},\mathbf{\Omega}_{L}}\hskip 5.69046ptS_{R}$ (31a)
$\hskip 42.67912pt\text{s.t.}\hskip 5.69046pt\operatorname{Tr}(\mathbf{WW}^{\mathrm{H}})\leq P_{T}$ (31b)
$\hskip 65.44142pt\gamma_{i}\leq\rho_{s}\hskip 8.5359pt\text{for}\hskip 5.69046pti\in\{L,E\}$ (31c)
$\hskip 65.44142pt|e^{j\varphi_{l}}|=1,\hskip 8.5359pt\text{for}\hskip 5.69046ptl\in\{1,\cdots,N_{L}\}$ (31d)
where $S_{R}=\sum\limits_{k=1}^{K}S_{R,k}$ is the achievable sum secrecy rate of legitimate users and $\rho_{s}$ is the maximum achievable sensing SINR rate. Here, since both legitimate and eavesdropper UAVs communicate directly with the BS without any reflected signals from RISs, the beamforming optimization can simply be independent of $L$-RIS configuration. This allows the problem (P1) to be decomposed into a two-stage optimization framework that first optimizing the transmit beamforming matrix $\mathbf{W}$, and later optimizing the $L$-RIS reflection matrix $\mathbf{\Omega}_{L}$.
III-A Beamforming Optimization
In the proposed scheme, in order to optimize the beamforming matrix $\mathbf{W}$ ensuring achievable sensing performance for both legitimate and eavesdropper UAV targets, the following optimization problem is formulated:
$\text{(P2)}\hskip 11.38092pt\max_{\mathbf{W}}\min\hskip 5.69046pt\gamma_{i}$ (32a)
$\hskip 56.9055pt{\text{s.t.}}\hskip 5.69046pt\gamma_{i}\leq\rho_{s}$ (32b)
$\hskip 76.82234pt\operatorname{Tr}(\mathbf{WW}^{\mathrm{H}})\leq P_{T}$ (32c)
Here, considering (8-9), for $\mathbf{A}_{i}\in\mathbb{C}^{T_{x}\times T_{x}}=\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})$ being the round trip channel between BS and $i$-UAV target, the sensing SINR expression in (9) can be re-expressed as
$\gamma_{i}=\frac{\operatorname{Tr}(\mathbf{A}_{i}^{\mathrm{H}}\mathbf{A}_{i}\mathbf{R_{x}})}{\operatorname{Tr}(\mathbf{A}_{j}^{\mathrm{H}}\mathbf{A}_{j}\mathbf{R_{x}})+\sigma_{n}^{2}}$ (33)
where $j\neq i\in\{L,E\}$. Therefore, using a semi-definite relaxation (SDR)-based approach, (P2) can be reformulated as
$\text{(P2)}\hskip 11.38092pt\max_{\mathbf{R_{x}}}\hskip 2.84544pt\min\hskip 5.69046pt\gamma_{i}$ (34a)
$\hskip 54.06006pt{\text{s.t.}}\hskip 8.5359pt\gamma_{i}\leq\rho_{s}$ (34b)
$\hskip 73.97733pt\operatorname{Tr}(\mathbf{A}_{i}^{\mathrm{H}}\mathbf{A}_{i}\mathbf{R_{x}})\geq\operatorname{Tr}(\mathbf{A}_{j}^{\mathrm{H}}\mathbf{A}_{j}\mathbf{R_{x}})+\sigma_{n}^{2}$ (34c)
$\hskip 73.97733pt\operatorname{Tr}(\mathbf{R_{x}})\leq P_{T}$ (34d)
using the SeDuMi solver within the CVX optimization toolbox [29]. Subsequently, the beamforming matrix $\mathbf{W}$ can be directly obtained through eigenvalue decomposition (EVD) [17].
III-B Legitimate RIS Phase Optimization
In this subsection, after the beamforming matrix $\mathbf{W}$ is obtained, the reflection coefficients of $L$-RIS are optimized to maximize the secrecy rate of the $K$ legitimate users. From (30), since the secrecy rate $S_{R}$ is the total sum of the achievable secrecy rates across all legitimate users, and eavesdropper SINR of $E$-UAV target (29) across users are independent from $L$-RIS reflection, the $\boldsymbol{\Omega}_{L}$ (1) can be optimized to enhance $S_{R}$ by maximizing sum of achievable rates of all users. Therefore, the optimization of $\mathbf{\Omega}_{L}$ in (P1) can be converted to following problem:
$\text{(P3)}\hskip 11.38092pt\max_{\mathbf{\Omega}_{L}}\hskip 2.84544pt\min\eta_{k}$ (35a)
$\hskip 36.98866pt\text{s.t.}\hskip 5.69046pt|e^{j\varphi_{l}}|=1\hskip 5.69046pt\text{for}\hskip 5.69046ptl\in\{1,\cdots,N_{L}\}.$ (35b)
However, due to its non-concave objective function and non-convex unit-modulus constraints, the problem (P3) is difficult to solve [30]. Therefore, an SDR-based approach is adopted to simplify (35).
As a first step, the communication SINR given in (27) is re-expressed in a quadratic form by applying appropriate mathematical modifications. Here, applying trace equality [17], the numerator in (27) can be rewritten as $\left\|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{k}\right\|_{\mathrm{F}}^{2}=\operatorname{Tr}(\mathbf{\bar{H}}_{L,k}\mathbf{\Omega}_{L}^{\mathrm{H}}\mathbf{\bar{G}}_{L,k}^{\mathrm{T}}\mathbf{\Omega}_{L})=\mathbf{z}_{L}(\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,k})\mathbf{z}_{L}^{\mathrm{H}}$, where $\mathbf{\bar{H}}_{L,k}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{H}_{L}\mathbf{w}_{k}\mathbf{w}_{k}^{\mathrm{H}}\mathbf{H}_{L}^{\mathrm{H}}$ and $\mathbf{\bar{G}}_{L,k}^{\mathrm{T}}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{g}_{L,k}^{\mathrm{H}}\mathbf{g}_{L,k}$, while $\mathbf{z}_{L}\in\mathbb{C}^{1\times N_{L}}$ is the reflection vector of $L$-RIS that composes non-zero diagonal elements of reflection matrix $\mathbf{\Omega}_{L}$. In a similar way, the denominator of $\eta_{k}$ can be rewritten in quadratic form as $\left\|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{\hat{k}}\right\|_{\mathrm{F}}^{2}=\mathbf{z}_{L}(\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,\hat{k}})\mathbf{z}_{L}^{\mathrm{H}}$ and $\left\|\mathbf{g}_{M,k}\mathbf{\Omega}_{M}\mathbf{H}_{M}\mathbf{W}\right\|_{\mathrm{F}}^{2}=\mathbf{z}_{M}(\mathbf{\bar{H}}_{M,k}\odot\mathbf{\bar{G}}_{M,{k}})\mathbf{z}_{M}^{\mathrm{H}}$, where $\mathbf{\bar{G}}_{L,\hat{k}}^{\mathrm{T}}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{g}_{L,\hat{k}}^{\mathrm{H}}\mathbf{g}_{L,\hat{k}}$, $\mathbf{\bar{H}}_{M,k}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathbf{H}_{M}\mathbf{R_{x}}\mathbf{H}_{M}^{\mathrm{H}}$ and $\mathbf{\bar{G}}_{M,k}^{\mathrm{T}}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathbf{g}_{M,k}^{\mathrm{H}}\mathbf{g}_{M,k}$, while $\mathbf{z}_{M}\in\mathbb{C}^{1\times N_{M}}$ is the random interference vector of $M$-RIS that includes non-zero diagonal elements of malicious reflecting matrix $\mathbf{\Omega}_{M}$. Therefore, (27) can be rewritten as
$\eta_{k}=\frac{\operatorname{Tr}(\mathbf{C}_{L,k}\mathbf{Z}_{L})}{\sum_{\hat{k}}\operatorname{Tr}(\mathbf{C}_{L,\hat{k}}\mathbf{Z}_{L})+\operatorname{Tr}(\mathbf{C}_{M,\hat{k}}\mathbf{Z}_{M})+\sigma_{n}^{2}}$ (36)
where $\mathbf{C}_{L,k}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,k}$, $\mathbf{C}_{L,\hat{k}}=\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,\hat{k}}$ and $\mathbf{C}_{M,k}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathbf{\bar{H}}_{M,k}\odot\mathbf{\bar{G}}_{M,k}$, while $\mathbf{Z}_{L}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{z}_{L}^{\mathrm{H}}\mathbf{z}_{L}$ and $\mathbf{Z}_{M}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathbf{z}_{M}^{\mathrm{H}}\mathbf{z}_{M}$. Therefore, the legitimate phase optimization problem in (35) can be converted to the following quadratically constrained quadratic programming (QCQP) problem as follows
$\text{(P3)}\max_{\mathbf{Z}_{L}}\hskip 2.84544pt\min\eta_{k}$ (37a)
s.t.
$\operatorname{Tr}(\mathbf{C}_{L,k}\mathbf{Z}_{L})-\Big{(}\sum_{\hat{k}}\operatorname{Tr}(\mathbf{C}_{L,\hat{k}}\mathbf{Z}_{L})+\operatorname{Tr}(\mathbf{C}_{M,\hat{k}}\mathbf{Z}_{M})+\sigma_{n}^{2}\Big{)}\geq 0$
$\hskip 39.83368pt\mathbf{Z}_{L}\succeq 0$ (37c)
$\hskip 39.83368pt\mathbf{Z}_{L}(l,l)=1\hskip 5.69046pt\text{for}\hskip 5.69046ptl\in\{1,\cdots,N_{L}\}.$ (37d)
After relaxing the non-convex constraint (37d) via an SDR-based approach, (P3) can be efficiently solved using the CVX optimization toolbox [29]. However, since the obtained solution is not always guaranteed to be rank-one, additional Gaussian approximation or EVD can be applied to get a feasible rank-one solution [31].
Overall, for $\varepsilon>0$ being the solution accuracy, the complexity of SDR-based solution for beamforming optimization problem (P2) in (34) is calculated as $\mathcal{O}(T_{x}^{4.5}\log(1/\varepsilon)$ while for legitimate phase optimization problem (P3) in (37), it is $\mathcal{O}(N_{L}^{4.5}\log(1/\varepsilon))$ [31].
IV Performance Evaluation
In this section, the secrecy rate and sensing performance of the proposed UAV-mounted RIS-assisted secure ISAC system are evaluated through Monte Carlo simulations under various scenarios. A BS with $T_{x}=8$ transmit antennas, $K=4$ legitimate users and two UAV sensing targets, where one is equipped with a legitimate RIS and the other with a malicious RIS, are considered. The BS is positioned at a height of $5$ m above ground level, and both UAV targets are deployed at $10$ m height. The users are randomly distributed at ground level, where the details of the placements are illustrated in Fig. 2. Additionally, the simulation parameters are summarized in Table I.
Table I: Simulation Parameters
Parameter Value Description
$f_{c}$ $3.5$ GHz Carrier frequency
$T_{x}$ 88 Number of transmit antennas
$K$ 44 Number of legitimate users
$M$ 22 Number of UAV sensing targets
$L_{s}$ $100$ Coherent time length
$\kappa$ 44 dB Rician factor
$\alpha_{n}$, $\alpha_{k}$ $2.2$ Path loss exponent
$S$ 11 $\text{m}^{2}$ RCS
$L_{0}$ 30-30 dB Reference path attenuation at $1$ m
$\sigma^{2}_{n}$ 120-120 dBW Noise power
$\theta^{\text{BS}}_{L},\phi^{\text{BS}}_{L}$ $(-141^{\circ},-9^{\circ})$ AoD pairs for $L$-UAV target
$\theta^{\text{BS}}_{E},\phi^{\text{BS}}_{E}$ $(-160^{\circ},-5^{\circ})$ AoD pairs for $E$-UAV target
$d_{L}$ $32.4$ m Distance from BS to $L$-UAV
$d_{E}$ $58.7$ m Distance from BS to $E$-UAV
$d_{L,k}$} $\{11.3,10.6,10.4,10.8\}$ m Distance from $L$-RIS to $\text{U}_{k}$}, for $k\in\{1:K\}$
$d_{M,k}$ $\{36.4,34.5,33.6,32.8\}$ m Distance from $M$-RIS to Uk, for $k\in\{1:K\}$
Figure 2: Spatial layout of the simulation environment for the BS, the RISs, and the users.
Fig. 3 presents the sum secrecy rate of four users, $S_{R}$, as a function of the total transmit power of the BS, $P_{T}$, for $N_{L}=144$ with varying numbers of malicious RIS elements, $N_{M}\in\{49,144,256\}$. These results illustrate the impact of the malicious RIS attacks on $S_{R}$ as $P_{T}$ increases, such that a larger malicious RIS size $N_{M}$ generates stronger interference on the legitimate communication links, and hence $S_{R}$ degrades. Furthermore, in the presence of the malicious RIS (i.e., $N_{M}\in\{49,144,256\}$), it is observed that when $P_{T}$ increases beyond a certain point ($P_{T}=6$ dBW), the $S_{R}$ performance starts degrading. These results can be attributed to the fact that, although a higher $P_{T}$ improves the achievable rate of the legitimate users, it simultaneously increases the sensing interference experienced by the users and enhances the SINR of the $E$-UAV target. This highlights the critical trade-off between enhancing legitimate user rates and mitigating interference effects, which necessitates the appropriate selection of $P_{T}$ values.
Figure 3: Comparison of the sum secrecy rate of the proposed scheme with reference [24] for varying $N_{M}$ sizes.
Fig. 3 also compares the $S_{R}$ performance of the proposed UAV-mounted RISs-assisted ISAC scheme with that of the benchmark RIS-aided ISAC systems in [24], where one of the targets acts as an eavesdropper without involving a malicious RIS. The results indicate that the proposed system under dual security threats of eavesdropper target and malicious RIS attacks, and the benchmark scheme affected only by an eavesdropper target [24] exhibit a similar $S_{R}$ behavior. As it is expected, the benchmark scheme [24] achieves better $S_{R}$ performance, and performance degradation after a certain point occurs at a higher transmit power ($P_{T}=10$ dBW). This result indicates that the benchmark scheme [24], which operates under a single target eavesdropping security threat, is naturally more robust than the proposed scheme that tackles the more challenging dual-threat scenario of simultaneous eavesdropping and malicious attacks.
Figure 4: Comparison of sum secrecy rate of the proposed scheme for varying $N_{L}$ sizes.
Figure 5: Sum secrecy rate performance of the proposed scheme for increasing $N_{M}$ sizes.
Figure 6: Sum secrecy rate performance of the proposed scheme for varying $P_{T}$ values.
In Fig. 4, the $S_{R}$ performance of the proposed scheme with a malicious RIS size of $N_{M}=36$ is evaluated for varying $N_{L}$. The results demonstrate that when $N_{L}$ is significantly larger than $N_{M}$ ($N_{L}=100,121,144$), the $S_{R}$ improves with increasing $P_{T}$, as the abundance of legitimate links offsets the impact of malicious links. However, when the size of $N_{L}$ equals $N_{M}$ ($N_{L}=36$), $S_{R}$ follows an irregular pattern as $P_{T}$ increases, while still maintaining secure communication. This irregularity arises from the comparable strengths of the legitimate and malicious reflections. It can be deduced from the results that $N_{L}$ must be sufficiently larger $N_{M}$ to effectively counteract malicious RIS attacks.
Fig. 5 presents $S_{R}$ performance of the proposed system as a function of $N_{M}\in\{25\sim 225\}$, with $N_{L}=100$ legitimate RIS and varying $P_{T}\in\{-3,0,3,6\}$ dBW. As expected, the results demonstrate that increasing $P_{T}$ improves $S_{R}$ performance. Conversely, a larger $N_{M}$ size intensifies the random interference attacks, which degrade its performance. However, the results also reveal that the proposed scheme sustains secure communication even at relatively low $P_{T}$ values and large $N_{M}$, demonstrating the effectiveness of the proposed algorithm.
Figure 7: Sensing SINR of both UAV targets for varying horizontal distance $d^{x}_{E}$.
Figure 8: Root CRB of the $E$-UAV for varying horizontal distance $d^{x}_{E}$.
In Fig. 6, the $S_{R}$ performance of the proposed scheme for increasing $N_{L}$ sizes is evaluated. It can be clearly observed from the Fig. 6 that an increase in $N_{L}$ enhances the $S_{R}$ performance of the proposed scheme, as it establishes a greater set of legitimate links for communication users. The figure also clearly demonstrates that $S_{R}$ improves as $P_{T}$ increases.
In the following, the sensing performance of the proposed system is evaluated in terms of sensing SINR of both the legitimate and the eavesdropper targets, and the root CRB estimations for AoDs of the eavesdropper UAV target, as illustrated in Figs. 7 and 8, respectively.
In Fig. 7, to evaluate the effectiveness of the proposed algorithm on the sensing SINR performance of UAV targets, the SINR values for both the legitimate and eavesdropper UAVs as a function of $P_{T}$ across varying horizontal distances ($d_{E}^{x}\in 25\sim 50$) of the $E$-UAV are presented. As it can be clear from Fig. 7(a) and 7(b), at a distance of $d_{E}^{x}=25$ m, the UAV targets become at the same distance from the BS (see Table I), yielding similar sensing SINR values for both targets. However, as illustrated in Fig. 7(a), when the $E$-UAV moves farther away, its sensing SINR decreases due to higher path loss, whereas the $L$-UAV’s sensing SINR improves as it experiences less interference from the $E$-UAV.
To provide further insights into sensing performance evaluation, in Fig. 8, the root CRB estimations for AoDs of the $E$-UAV are provided across varying horizontal distances of $E$-UAV, $d_{E}^{x}\in 25\sim 55$ m. As expected, a larger $d_{E}^{x}$ degrades root CRB estimation due to weakened sensing signal strength caused by higher path attenuation. It is also apparent from the results that an increase in $P_{T}$ improves the root CRB estimations.
V Conclusion
In this study, a target-mounted RISs-aided ISAC framework has been proposed to ensure secure communication under dual security threats: an eavesdropper target that intercepts legitimate communications and a malicious RIS that launches random interference attacks. To address these challenges, a non-convex optimization problem is formulated to maximize the sum secrecy rate of the overall system. Then, in order to solve this problem, an SDR-based two-stage algorithm has been developed to optimize the beamforming matrix and phase shifts of the legitimate RIS. Furthermore, comprehensive computer simulations have been conducted to investigate the effectiveness of the proposed algorithm on the secrecy rate and sensing performance metrics across various system configurations. Future research will focus on leveraging RISs to strengthen the security of sensing signals, particularly in dynamic scenarios involving moving targets.